OpenStudy (anonymous):

What is m/-6 + 6 = 41?

OpenStudy (anonymous):

Is it (m/-6) + 6 = 41 ?

OpenStudy (anonymous):

Yes! Sorry, I should have typed it better.. Lol

OpenStudy (anonymous):

So, subtract 6 from both sides: (m/-6) + 6 = 41 (m/-6) + 6 - 6 = 41 - 6 (m/-6) = 35 Then multiply both sides by (-6) (m/-6) (-6) = 35(-6) m = -210

OpenStudy (anonymous):

Does that make sense? You just have to move the numbers around to solve for m, but it might take a couple of steps

OpenStudy (anonymous):

Another way to solve...:D \[\frac{ m }{ -6 }+6=41\] \[\frac{ m-36 }{-6 }=41\] \[-6*\frac{ m-36 }{ -6 }=41*-6\] \[m-36=-246\] \[m-36+36=-246+36\] \[m=-210\]